# Unsupervised detection of semantic correlations in big data

### v2

@@ III Comparison with other methods
First, for a small 30x30 ferromagnetic Ising model we studied the thermal dependence of the ID (Fig. 4, left). At low temperature the system is in the ferromagnetic phase, which implies that a large fraction of the spins points in the same direction. We observe for all estimators except FCI that in these conditions the estimated ID is small and of roughly comparable magnitude. The estimates are consistent because in this special regime the local estimators (MLE, GRIDE, I3D and CD) are not severely affected by the dimensionality. In the same regime, the FCI estimator predicts a qualitatively different behavior, with the ID being very high and a decreasing function of T. This is at odds with physical intuition. Indeed, in a Ising model at low temperature only few defects are present, and the state of the system can be fully described by providing only the position of those defects. Therefore, the ID should smoothly tend to zero for $T\rightarrow 0$, similarly to the entropy. We notice that in the ferromagnetic phase most of the spins have the same value. In these conditions the centering and normalization procedure at the basis of FCI places the data on a hypersphere, but the distribution of the data on it is highly non-uniform. The qualitatively different results obtained by FCI might then be induced by the violation of one of its assumptions, that the data are isotropic when observed from the geometric center of the data set. In the paramagnetic phase, the BID estimate is instead comparable to the FCI estimate. Indeed, in this phase the spins are not polarized, namely the magnetization defined as $M=\sum_{i=1}^{N}\sigma_{i}$ is null. In these conditions, after centering and normalization, the data will be approximately isotropic, making the estimate by FCI more reliable. In contrast, all the other local estimators subestimate the value of the ID because they need an exponential number of samples in order to converge.
@@
The right panel of Fig. 4 shows the size scaling of the different algorithms at fixed low temperature, where we know the system is extensive, and thus the ID must scale linearly with the number of bits, $N$. For small system sizes we observe that the local estimators (MLE, GRIDE, I3D and CD) give results comparable with the BID, but their estimate of the ID does not scale linearly with $N$, at odds with physical evidence. This is the behavior which prompted us to develop a new estimator, specifically designed for bits or spins. The BID scales linearly with the number of bits, and so does FCI. Nonetheless for FCI the slope of the line is one order of magnitude bigger and, as discussed above, the corresponding estimation is also at odds with physical intuition.
@@
We finally performed ID estimates with GRIDE and FCI on the images and text databases. For our experiment on images, we observe that both GRIDE and FCI behave qualitatively similarly to the BID, namely the plateau observed for real data and its absence for images constructed grouping random patches are reproduced (see Sec. A.8.1 of Supp. Inf.). However, we remark that the ID estimation performed with the GRIDE is more than one order of magnitude smaller than the estimations done by BID, whereas the estimations of FCI are systematically below that of BID, by roughly a factor of 2. Additionally, in Sec. A.8.3 of Supp. Inf. we study the BID of simple images directly in pixel space, for MNIST and CIFAR10. There we show that the BID of MNIST is fully compatible with other real-space estimators, and we discuss the challenges of finding correct binarizations for RGB images. For the text experiment, we show in Sec. A.8.2 of Supp. Inf. that GRIDE fails to display a plateau at the semantic boundary since, as we showed in Fig. 4, it doesn’t scale correctly with system size. Instead, the FCI estimator produces qualitatively comparable results to ours, reproducing the linear scaling for the ID at the layer zero and detecting the semantic boundary in the case of concatenated paragraphs. Nonetheless, we observed significant differences in the size scaling exponents of the BID normalized per bit at the last layer of the transformer, namely the slope $\eta$ of the lines in the log-log plot of $BID(T)/N\propto T^{-\eta}$. Whereas for the BID we get an exponent of $\eta_{BID}=0.42$, for FCI and GRIDE we get much higher values, $\eta_{FCI}=0.89$, and $\eta_{GRIDE}=0.70$. The study of the relationship between the BID and the scaling exponents in systems with power-law correlations, and in particular in critical systems, together with the quantitative discrepancies observed, are interesting topics which we are currently investigating.
@@ IV Conclusions
It is noteworthy that the last two datasets consist of real-valued features, while the BID estimator is designed for binary data. Each feature was converted into a binary variable by taking its sign. In Supp. Inf. we show that the trends observed in Fig. 3 hold when using two bits per feature instead of one. Specifically, the relative BID estimated for this higher-precision representation scales with the same power law. This suggests that when the intrinsic dimension becomes extremely large (around 1000 or more), its trends are insensitive to the precision used to represent individual features, and binning retains the essential information of the representation, as suggested in Refs. Hubara et al. [2018, 2016].
@@
The code used to obtain the results of this work can be found in https://github.com/acevedo-s/BID. Routines to compute the BID are implemented in the open-source package DADApyGlielmo et al. [2022b].
@@ Acknowledgements
The authors acknowledge Marco Baroni, Antonello Scardicchio and Marcello Dalmonte for useful discussions.
Author contributions
SA performed the numerical experiments. SA, AR, and AL designed the experiments, analyzed the data, and wrote the manuscript.
Competing interests
The authors declare no competing interests.
Appendix A Supplementary Information
A.1 Optimization
The meta-parameter $r^{*}$ is computed as the quantile of order $\alpha^{*}$ of the empirical probability of distances $P_{emp}(r)$, i.e., $\alpha^{*}=\sum_{r=0}^{r^{*}}P_{emp}(r)$, and we typically select the value of $\alpha^{*}$ that gives the lowest KL divergence (4). The minimization of (4) is performed stochastically, meaning we propose small random moves from $(d_{0},d_{1})$ to $(d_{0}+\delta_{0},d_{1}+\delta_{1})$ for suitable $\delta_{0}$ and $\delta_{1}$ and we take them if they decrease the loss function, or reject them otherwise. When comparing BID’s of a system computed with different model parameters (e.g., temperature, or size) we always do it at a shared value of $\alpha^{*}$. Occasionally the smallest sampled distances are noisy and poorly sampled, which may induce overflow in the numerical minimization of (4), especially for very large bit streams. In this cases we introduce a regularizer $r_{min}$ defined as the quantile of order $\alpha_{min}$ of $P_{emp}$ to remove distances with $r<r_{min}$. Table 2 summarizes the values of these meta-parameters used in this work.
As an example we show additional optimization results for the 2D Ising model on the square lattice (Fig 1A,D,G from main text). Fig. 5 shows the thermal dependence of $d_{1}$, $\log{(KL)}$ and $r^{*}$ for different system sizes $L$. Notably $d_{1}$ has a maximum around $T_{c}$. Even if the error is maximum at $T_{c}$, the maximum of $\log{(KL)}$ is close to $10^{-7}$, implying that the model fits accurately the data, as shown in Fig. 1G. $r^{*}$ instead develops an inflexion point around $T_{c}$.
Figure 5: Thermal dependence of $d_{1}$, $\log{(KL)}$ and $r^{*}$ for the 2D Ising model on the square lattice, $N=L^{2}$ stands for the total number of spins. The dashed line corresponds to the exact critical temperature in the thermodynamic limit.
Figure / Table $\alpha_{min}$ $\alpha_{max}$
Table 1 0 1
Fig. 1A,D,G,B,E,H 0 .1
Fig. 1C,F,I 0 .05
Fig. 2 $10^{-2}$ 0.5
Fig. 3,9,8 $10^{-5}$7
Fig. 6 0 0.4
Fig. 7 0 0.4
Table 2: Meta-parameters used for figures and tables.
A.2 Dependence on the number of samples
Fig. 6 shows the dependence of the BID on the number of samples, $N_{s}$, for the 2D ferromagnetic Ising model. The algorithm provides a stable estimate after a few hundred samples even for high ID values because the fit of the model is performed on a global observable, the probability of observing Hamming distance r in the dataset, $P_{emp}(r)$, a quantity which is well defined even in the presence of measure concentration in high dimensions. The right panel of Fig. 6 shows that the smallest sampled scales are those strongly affected by the size of the dataset, but larger values of $r$ are stable, i.e., the histograms computed for the different values of $N_{s}$ overlap. Note that Eq. (4) weights the contribution of the discrepancies with $P_{emp}(r)$, and thus the most statistically relevant distances dominate the optimization.
Figure 6: Left panel: BID normalized per bit versus the number of samples, for different temperatures, $T$. Right panel: Example histograms for $T=3$, and three different values of $N_{s}$. In solid black lines, the fits to each histogram (they overlap).
A.3 Bayesian inference of the BID by Markov Chain Monte Carlo
Figure 7: Empirical posterior density for the two parameters of our model. $d_{0}^{*}/N=0.0512$ and $d_{1}^{*}=1.9079$ correspond to the solution found optimizing Eq. (2). The posterior means and standard deviations are $\mu_{d_{0}}/N=0.0510$, $\sigma_{d_{0}}/N=0.0006$, $\mu_{d_{1}}=1.908$, $\sigma_{d_{1}}=0.002$ Number of samples: $N_{s}=500$, number of spins $N=10^{4}$, temperature $T=2.3.$
The parameters of model (2) can be found minimizing Eq. (4), as we did in our work, or alternative using the Metropolis Hastings algorithm on model (2), with a flat prior. Here we sample the probability distribution
@@ A.6 The BID of text representations
Figure 8: BID per bit calculated using all token representations in the sequence as a function of the number of tokens, for OPT350m on OpenWebtext. l = 0 stands for the embedding layer, l = 24 corresponds to the last transformer layer before the head. "Concat" stands for phrases constructed concatenating 150 tokens from two randomly selected data samples. The dashed line corresponds to the semantic boundary between the two phrases.
@@
Figure 9: Comparing BID size scalings using direct sign binarization (Nbits=1) or the binarization given by (8).
A.7 Maximum Likelihood Estimation of a scale-dependent BID
@@
Figure 10: Scale-dependent BID estimated by MLE (12) as a function of the observational scale $r$, on the two-dimensional ferromagnetic Ising model on the square lattice, for several temperatures $T$. Both axis are normalized by the total number of spins $N$ in the system. The two panels are separated temperatures above (right) and below (left) the critical temperature of the system for visual clarity.
@@
A.8 Comparison with other methods
A.8.1 ID dependence with crop size in image classification
Fig. 11 shows the intrinsic dimension computed using GRIDE and FCI on the same effective spin system of Fig. 2 for an example relative depth of 0.39. For the GRIDE estimator we used the parameter $k=4$ to select the scale.
Figure 11: BID results from Fig 2 compared with FCI and GRIDE on the same spin dataset.
A.8.2 ID of Text representations
Fig. 12 shows the ID estimated by BID,FCI and GRIDE on the effective spin system of Fig. 3. The two curves computed at layer zero (embedding layer) with FCI are cut at 200 tokens since the implementation present in https://github.com/vittorioerba/pyFCI crashed due to overflow.
Figure 12: BID results from Fig 3 compared with FCI and GRIDE on the same spin dataset. For the last layer we have that $BID(T)/N\propto T^{-\eta}$. For the BID we get an exponent of $\eta_{BID}=0.42$. For FCI and GRIDE we get much higher values, $\eta_{FCI}=0.89$, and $\eta_{GRIDE}=0.70$. Number of samples used: $N_{s}=5000$.
A.8.3 BID of MNIST and CIFAR10
In this section we compute the BID of images directly on data space, for the MNIST and CIFAR10 datasets. For MNIST, different ID estimatorsDenti et al. [2022], Spigler et al. [2020] predict values between 10 and 15. The natural binary representation of MNIST is obtained rounding the floating-point values of the gray scale to either black or white. We found that this operation increases the value of the intrinsic dimension when computed with standard approaches. For example, following Ref. Spigler et al. [2020], the intrinsic dimension (there called effective dimension, $d_{eff}$) is obtained as the scaling exponent relating the average nearest-neighbor distance $\langle\delta\rangle$ with the number of samples $N_{s}$, namely $\langle\delta\rangle\sim N_{s}^{-\frac{1}{d_{eff}}}$. Through a linear fit with $N_{s}\in[4000,8000]$, we get $d_{eff}(MNIST)\approx 16$ and $d_{eff}(round(MNIST))\approx 22$. The value obtained by our estimator is consistent with those numbers, as we found $BID(MNIST)\approx 19$.
In the case of colored images it is no longer trivial to find a meaningful binarization that approximately preserves the information. The rounding procedure used for MNIST applied on each channel of the RGB images of CIFAR10 gives a BID of the order of $200$, at odds with reported valuesPope et al. [2021], Denti et al. [2022], Spigler et al. [2020] ranging between 15 and 35. This discrepancy is not surprising since our estimator is specifically designed to work with binary systems, and by binarizing complex RGB images one losses a lot of information. In our main manuscript, together with Sec. A.8.1, we showed that the representations of natural images inside deep neural networks allow to be meaningfully binarized, at least with respect to the measurement of the intrinsic dimension. The construction of meaningful binary representations of colored images and their impact on the corresponding BID remains a very interesting open line of research.
A.9 Neighbourhoods are approximately preserved after binarizing high dimensional activations
In this section we study how the binarization of activations through the sign function affects the neighbourhood structures inside a LLM. To do this we start by considering the nearest neighbour of each data point in the full-precision space of activations generated by Pythia 410m on 6000 samples of Wikitext with 300 tokens each. For every such first neighbour we then compute its rank in the system of binarized activations. While in continuous space distances do not repeat and each rank is unique, for discrete systems this is no longer true and one has to choose a definition for the rank that takes this into account. In the following we consider that equidistant points share the rank.
Fig. 13 shows the distribution of ranks obtained after binarizing next-neighbour vectors of activations. For both initial (0) and last (24) layers the distributions are peaked at $1$, meaning that the most probable rank after binarization is still 1 (i.e. first neighbours). Moreover, the expected rank after binarizing remains smaller than 10 in both cases, proving that ranks are approximately preserved after binarization. For the (unnormalized) initial layer we observed that substracting the layer-mean was beneficial to concentrate the post-binarization rank distribution closer to 1.
Figure 13: Distribution of ranks obtained after binarizing the activations of a LLM, using points that were first neighbours in the full precision space. On vertical dashed lines the expected value of the distribution.

### v3

@@ Results
First, for a small 30x30 ferromagnetic Ising model we studied the thermal dependence of the ID (Fig. 4-a)). At low temperature the system is in the ferromagnetic phase, which implies that a large fraction of the spins points in the same direction. We observe for all estimators except FCI that in these conditions the estimated ID is small and of roughly comparable magnitude. The estimates are consistent because in this special regime the local estimators (MLE, GRIDE, I3D and CD) are not severely affected by the dimensionality. In the same regime, the FCI estimator predicts a qualitatively different behavior, with the ID being very high and a decreasing function of T. This is at odds with physical intuition. Indeed, in a Ising model at low temperature only few defects are present, and the state of the system can be fully described by providing only the position of those defects. Therefore, the ID should smoothly tend to zero for $T\rightarrow 0$, similarly to the entropy. We notice that in the ferromagnetic phase most of the spins have the same value. In these conditions the centering and normalization procedure at the basis of FCI places the data on a hypersphere, but the distribution of the data on it is highly non-uniform. The qualitatively different results obtained by FCI might then be induced by the violation of one of its assumptions, that the data are isotropic when observed from the geometric center of the data set. In the paramagnetic phase, the BID estimate is instead comparable to the FCI estimate. Indeed, in this phase the spins are not polarized, namely the magnetization defined as $M=\sum_{i=1}^{N}\sigma_{i}$ is null. In these conditions, after centering and normalization, the data will be approximately isotropic, making the estimate by FCI more reliable. In contrast, all the other local estimators subestimate the value of the ID because they need an exponential number of samples in order to converge.
@@
Fig. 4-b) shows the size scaling of the different algorithms at fixed low temperature, where we know the system is extensive, and thus the ID must scale linearly with the number of bits, $N$. For small system sizes we observe that the local estimators (MLE, GRIDE, I3D and CD) give results comparable with the BID, but their estimate of the ID does not scale linearly with $N$, at odds with physical evidence. This is the behavior which prompted us to develop a new estimator, specifically designed for bits or spins. The BID scales linearly with the number of bits, and so does FCI. Nonetheless for FCI the slope of the line is one order of magnitude bigger and, as discussed above, the corresponding estimation is also at odds with physical intuition.
@@
We finally performed ID estimates with GRIDE and FCI on the images and text databases. For our experiment on images, we observe that both GRIDE and FCI behave qualitatively similarly to the BID, namely the plateau observed for real data and its absence for images constructed grouping random patches are reproduced (see Supplementary Note 3.1). However, we remark that the ID estimation performed with the GRIDE is more than one order of magnitude smaller than the estimations done by BID, whereas the estimations of FCI are systematically below that of BID, by roughly a factor of 2. Additionally, in Supplementary Note 3.3 we study the BID of simple images directly in pixel space, for MNIST and CIFAR10. There we show that the BID of MNIST is fully compatible with other real-space estimators, and we discuss the challenges of finding correct binarizations for RGB images. For the text experiment, we show in Supplementary Note 3.2 that GRIDE fails to display a plateau at the semantic boundary since, as we showed in Fig. 4, it doesn’t scale correctly with system size. Instead, the FCI estimator produces qualitatively comparable results to ours, reproducing the linear scaling for the ID at the layer zero and detecting the semantic boundary in the case of concatenated paragraphs. Nonetheless, we observed significant differences in the size scaling exponents of the BID normalized per bit at the last layer of the transformer, namely the slope $\eta$ of the lines in the log-log plot of $BID(T)/N\propto T^{-\eta}$. Whereas for the BID we get an exponent of $\eta_{BID}=0.42$, for FCI and GRIDE we get much higher values, $\eta_{FCI}=0.89$, and $\eta_{GRIDE}=0.70$. The study of the relationship between the BID and the scaling exponents in systems with power-law correlations, and in particular in critical systems, together with the quantitative discrepancies observed, are interesting topics which we are currently investigating.
@@ Conclusions
It is noteworthy that the last two datasets consist of real-valued features, while the BID estimator is designed for binary data. Each feature was converted into a binary variable by taking its sign. In Methods, section ‘The BID of text representations’ we show that the trends observed in Fig. 3 hold when using two bits per feature instead of one. Specifically, the relative BID estimated for this higher-precision representation scales with the same power law. This suggests that when the intrinsic dimension becomes extremely large (around 1000 or more), its trends are insensitive to the precision used to represent individual features, and binning retains the essential information of the representation, as suggested in Refs. Hubara et al. [2018, 2016].
@@
The code used to obtain the results of this work can be found in https://github.com/acevedo-s/BID. Routines to compute the BID are implemented in the open-source package DADApyGlielmo et al. [2022b]. A tutorial notebook can be found at https://github.com/sissa-data-science/DADApy/blob/main/examples/notebook_hamming.ipynb
@@ Acknowledgements
The authors acknowledge Marco Baroni, Antonello Scardicchio and Marcello Dalmonte for useful discussions. AL and SA acknowledge foundings from NextGenerationEU through the Italian National Centre for HPC, Big Data, and Quantum Computing (Grant No. CN00000013). AL and SA also acknowledge financial support by the region Friuli Venezia Giulia (project F53C22001770002)
Methods
Bayesian inference of the BID by Markov Chain Monte Carlo
Figure 5: BID Bayesian inference. Panels a) and b) show the empirical posterior densities for the two parameters of our model, $d_{0}$ and $d_{1}$, respectively. $d_{0}^{*}/N=0.0512$ and $d_{1}^{*}=1.9079$ correspond to the solution found optimizing Eq. (2). The posterior means and standard deviations are $\mu_{d_{0}}/N=0.0510$, $\sigma_{d_{0}}/N=0.0006$, $\mu_{d_{1}}=1.908$, $\sigma_{d_{1}}=0.002$. Number of samples: $N_{s}=500$, number of spins $N=10^{4}$, temperature $T=2.3.$
The parameters of model (2) can be found minimizing Eq. (4), as we did in our work, or alternative using the Metropolis Hastings algorithm on model (2), with a flat prior. Here we sample the probability distribution The parameters of model (2) can be found minimizing Eq. (4), as we did in our work, or alternative using the Metropolis Hastings algorithm on model (2), with a flat prior. Here we sample the probability distribution
@@ Methods
Figure 6: Stability of results against changes in both data and model. BID per bit calculated using all token representations in the sequence as a function of the number of tokens, for OPT350m on OpenWebtext. l = 0 stands for the embedding layer, l = 24 corresponds to the last transformer layer before the head. "Concat" stands for phrases constructed concatenating 150 tokens from two randomly selected data samples. The dashed line corresponds to the semantic boundary between the two phrases.
@@
Figure 7: Stability with respect to binarization protocol. Comparison of BID size scalings using direct sign binarization (Nbits=1) or the binarization given by (8).
Maximum Likelihood Estimation of a scale-dependent BID
@@
Figure 8: Maximum likelihood estimation of local BID. Scale-dependent BID estimated by MLE (12) per bit, $\hat{d}(r)/N,$ as a function of the observational scale $r$, on the two-dimensional ferromagnetic Ising model on the square lattice, for several temperatures $T$. Both axis are normalized by the total number of spins $N$ in the system. The two panels are separated temperatures above (right) and below (left) the critical temperature of the system for visual clarity.
@@
Supplementary Material
Supplementary Material is available in the published open access version, Ref. Acevedo et al. [2025].